OpenStudy (anonymous):

what is 1/3x + 1/3(2x - 15)= 3 1/2

OpenStudy (anonymous):

x = 13/2

OpenStudy (anonymous):

(1/3)*(2*x-15)+(1/3)*x = 3*(1/2) // - 3*(1/2) (1/3)*(2*x-15)+(1/3)*x-(3*(1/2)) = 0 (1/3)*(2*x-15)+(1/3)*x-3/2 = 0 1/3*(2*x-15)+x/3-3/2 = 0 (1/3*2*3*(2*x-15))/(2*3)+(2*x)/(2*3)+(-3*3)/(2*3) = 0 1/3*2*3*(2*x-15)+2*x-3*3 = 0 6*x-30-9 = 0 6*x-39 = 0 (6*x-39)/(2*3) = 0 (6*x-39)/(2*3) = 0 // * 2*3 6*x-39 = 0 6*x-39 = 0 // + 39 6*x = 39 // : 6 x = 39/6 x = 13/2

OpenStudy (anonymous):

hope this helped

OpenStudy (anonymous):

Thank you so much... :)